### v1

# 3d Conformal Field Theories via Fuzzy Sphere Algebra

Abstract
Fuzzy sphere models conjecturally realize 3d CFTs in small systems of spinful fermions, but why they work so well is still not fully understood. Their Hamiltonians are built from electron density operators projected to the lowest Landau Level. We analyze the algebra of the density modes and verify that it satisfies the Jacobi identity. The fuzzy sphere geometry admits two thermodynamic limits: a local planar limit yielding the fuzzy plane, and a commutative limit yielding an ordinary sphere. In the planar limit, high-angular-momentum modes recover the Girvin–MacDonald–Platzman algebra, whereas in the commutative limit the low-angular-momentum modes become semiclassical. We further find an explicit representation of the conformal algebra $so(3,2)$ in the minimal two-electron system and extend it to larger systems via an $so(3)$ equivariant coproduct. Because the coproduct splits one $so(3)$ representation into a tensor product, it is structurally mismatched with the thermodynamic limit of critical fuzzy sphere Hamiltonians.
1 Introduction
3d topological quantum field theories (TQFTs) and 2d conformal field theories (CFTs) are among the best examples of quantum field theories that are both mathematically rigorous and physically relevant. They are related by an edge-bulk correspondence that is realized by quantum Hall materials. 4d TQFTs are suggested to be related to topological insulators and superconductors, while the 3d Ising model is conjectured to be conformally invariant at the critical parameters. Therefore, a generalization to one dimension higher of the 3d TQFT—2d CFT correspondence would be both interesting and daunting.
One possible mathematical approach to 2d CFTs, in particular the minimal models, is via the anyonic chains Zini and Wang (2018). This generalized spin chain approach is based on finite-dimensional Hilbert space approximations, hence obviously mathematical, but subtleties arise in the limiting process. Extensive numerical simulation and mathematical verification of simple examples provide strong evidence that this anyonic chain approach to 2d CFTs, at least the minimal models, is viable Feiguin et al. (2007); Zini and Wang (2018). In this paper, we initiate the first step for an extension of a generalized spin chain approach to 3d CFTs via fuzzy sphere models.
There are two important classes of algebras in the anyonic chain approach to 2d CFTs: the finite-dimensional Temperley-Lieb-Jones (TLJ) algebras $TLJ_{n}(A)$, and the infinite-dimensional Virasoro algebras. Both algebras play essential roles in the solution and understanding of the 2d Ising model and CFTs. At certain special values of the Kauffman variable $A$, the quotients of the Temperley-Lieb (TL) algebras $TL_{n}(A)$ by the Jones-Wenzl projectors lead to $\mathbb{C}^{*}$-TLJ algebras $TLJ_{n}(A)$ generated by TL generators $\{e_{i}\}_{i=1}^{n-1}$ Wang (2010). These TL generators $\{e_{i}\}$ could be regarded as finite lattice version of the Virasoro generators $\{L_{m}\}$ in the sense given by the Koo-Saleur formulas via the Fourier transform Koo and Saleur (1994); Zini and Wang (2018). Ultimately, we would like to find analogues of the TLJ and Virasoro algebras from the fuzzy sphere algebra and the diffeomorphism group of $S^{2}\times S^{1}$.
One 2+1d generalization of the anyonic chain is the fusion surface model Inamura and Ohmori (2024); Eck and Fendley (2025). Since we have not yet had success using them for approximating 3d CFTs, we instead propose to treat the fuzzy sphere Hamiltonians Zhu et al. (2023) as generalized spin chains (finite-dimensional quantum lattice systems in which we do not assume microscopic locality). The fuzzy sphere model describes spinful electrons on the 2-sphere with a monopole at the origin, projected to the lowest Landau level (LLL). Remarkably, it reproduces scaling dimensions of the 3d Ising CFT with good precision already in small systems of only $\mathcal{O}(10)$ electrons.
The Landau-level–projected density operators that enter the fuzzy-sphere Hamiltonians generate an intriguing algebra that is of independent interest, and which we study in this paper. After briefly reviewing the construction of fuzzy sphere Hamiltonians in Section 3, we connect the density modes both to fuzzy spherical harmonics from noncommutative field theory Madore (1992); Grosse et al. (1996) and to the matrix Lie algebra $su(2N)$ using their realization as fermionic bilinears (Section 4.1). Treating the commutator as defining an abstract algebra, we show that the bracket obeys the Jacobi identity and identify commuting subalgebras in 4.2. We then analyze two natural thermodynamic limits of the algebra: a local planar limit to the fuzzy plane and a commutative limit to an ordinary sphere in Sec. 5. The commutative, semiclassical regime is expected to be the more relevant one for the low-energy spectrum of the critical fuzzy-sphere Hamiltonians.
In 2+1 dimensions, conformal transformations are described by the finite-dimensional Lie algebra $so(3,2)$, reviewed in Section 2. In Section 6, we identify an explicit $so(3,2)$ representation in the minimal two-electron fuzzy sphere model and discuss how it can be lifted to larger Hilbert spaces using an $so(3)$ equivariant coproduct. We emphasize, however, that this coproduct splits a single $so(3)$ representation into a tensor product, and therefore does not coincide with the thermodynamic limit relevant for the critical fuzzy sphere Hamiltonians, where one increases the size of a single irrep.
Our results set up the foundation for taking a low energy scaling limit of the fuzzy sphere models realizing 3d CFTs. Since such a scaling limit is difficult and subtle to take, we leave it to the future. Another long-term direction is to search for symmetry enhancements beyond global $SO(3,2)$, potentially connected to diffeomorphisms of $S^{2}\times S^{1}$ – the 2-fold cover of the conformal compactification of $\mathbb{R}^{2,1}$. But in this paper, we will focus only on $SO(3,2)$.
2 3d CFTs
Above two dimensions, the conformal transformations are always global and only form finite-dimensional Lie groups. For Euclidean spacetime of dimension $n$, the conformal group is $SO(n+1,1)$, while for the Lorentzian signature (n1,1)(n-1,1), the conformal group is $SO(n,2)$. We are interested in $n=3$ and Minkowski spacetime (as the fuzzy sphere models are formulated as 2+1d quantum Hamiltonians), for which the relevant conformal algebra is $so(3,2)$.
2.1 Conformal invariance and covariance
A 3d (quantum) conformal field theory (CFT) is a quantum field theory with symmetry extending the Poincare group to one that includes scale and Möbius transformations. While the correlation functions of a CFT are invariant, the fields and states only transform covariantly under the action of the extended symmetry group. Therefore, representation theory of the extended symmetry group could be a powerful tool for constructing mathematically nontrivial CFTs, especially when the symmetry is infinite-dimensional, as in two dimensions.
2.2 Conformal $so(3,2)$ algebra
The $so(3,2)$ algebra has 10 antisymmetric generators $M_{AB}=-M_{BA}$ with $A,B\in\{1,2,3,4,5\}$. Two generators commute if all indices are distinct, $M_{AB},M_{CD}]=0$. The nontrivial commutation relations are
$[M_{AB},M_{AC}]=\begin{cases}iM_{BC}&\text{ if }A=4,5\\ -iM_{BC}&\text{ if }A=1,2,3\\ \end{cases}.$
In particular, the $so(3)$ rotation subalgebra is generated by
$J_{z}=M_{12},\qquad J_{\pm}=M_{23}\pm iM_{13}.$
A standard physics basis for the remaining generators is given by the dilatation $D$, translations $P_{A}$ and special conformal transformations $K_{A}$ ($A=1,2,3$):
$D=M_{45},\qquad P_{A}=M_{4A}+iM_{5A},\qquad K_{A}=M_{4A}-iM_{5A}.$ (1)
Following Celeghini and Del Olmo (2013), it is sometimes convenient to instead use two commuting Cartans and four pairs of ladder operators as generators of $so(3,2)$. We denote
$J_{z}$ $=M_{12},\qquad J_{\pm}=M_{23}\pm iM_{13},\qquad D_{z}=M_{45},\qquad D_{\pm}=M_{43}\pm iM_{53},$
where $D_{z}$ coincides with the dilatation operator $D$ and $D_{\pm}$ correspond to $K_{z}$ and $P_{z}$ in the usual physics convention (1)). The remaining four ladder operators are defined by commutators,
$R_{\pm}=\pm[J_{\pm},D_{\pm}],\qquad S_{\pm}=\pm[J_{\mp},D_{\pm}].$
In this basis, the commutation relations are given by
$=\pm J_{\pm},\qquad[J_{+},J_{-}]=2J_{z},\qquad[D_{z},D_{\pm}]=\pm D_{\pm},\qquad[D_{+},D_{-}]=-2D_{z},$ (2)
$[D_{z},Jz]$ $=0,\qquad[D_{z},J_{\pm}]=0,\qquad[D_{\pm},J_{z}]=0$
$[D_{\mp},R_{\pm}]$ $=\pm 2J_{\pm},\qquad[J_{\pm},S_{\pm}]=\pm 2D_{\pm}\qquad\text{ for }R_{\pm}=\pm[J_{\pm},D_{\pm}],\;S_{\pm}=\pm[J_{\mp},D_{\pm}],$
$[D_{\pm},R_{\pm}]$ $=0,\qquad[J_{\pm},R_{\pm}]=0,\qquad[D_{\pm},S_{\pm}]=0,\qquad[J_{\mp},S_{\pm}]=0.$
All remaining commutators follow from the ones above and the Jacobi identity.
2.3 Representations of $so(3,2)$
Because $so(3,2)$ is a noncompact Lie algebra, its finite-dimensional representations are all non-unitary. Unitary CFTs instead use infinite-dimensional, positive-energy unitary irreducible representations; an example of such is given by Dirac in Dirac (1963).
For a critical quantum Hamiltonian on the spatial sphere, emergent conformal symmetry is visible in the low-energy spectrum. In radial quantization, time translations are generated by the dilatation operator $D$, so scaling dimensions $\Delta$ are the eigenvalues of $D$. Via the state–operator correspondence, low-lying energy eigenstates correspond to local CFT operators labeled by $(\Delta,l)$: $l$ is the $so(3)$ angular momentum, and the energy is proportional to $\Delta$.
A useful way to understand this unitarity obstruction is to recall that the real Lie algebras $so(p,q)$ and $so(p+q)$ become isomorphic as complex Lie algebras after complexification, $so(p,q)_{\mathbb{C}}\cong so(p+q)_{\mathbb{C}}$. As a result, the same finite-dimensional complex representation can be viewed either as a representation of the non-compact real form $so(p,q)$ or the compact real form $so(p+q)$. While it is necessarily non-unitary for $so(p,q)$, it can admit a positive-definite invariant inner product—hence become unitary—when interpreted as a representation of $so(p+q)$. We illustrate this point below using the pair $so(2,1)$ and $so(3)$.
Both Lie algebras are generated by $\{L_{i}\}_{i=1}^{3}$ with Lie brackets
$L_{1},L_{2}]=i\sigma L_{3},\qquad[L_{2},L_{3}]=iL_{1},\qquad[L_{3},L_{1}]=iL_{2},$
with $\sigma=+1$ for $so(3)$ and $\sigma=-1$ for $so(2,1)$. Set $H=L_{3},E_{\pm}=L_{1}\pm iL_{2}$, then the Casimir is
$C=\sigma E_{\mp}E_{\pm}+H(H\pm 1).$
[…]
For $\sigma=+1$, the right hand side eventually becomes negative as $m$ gets too large in magnitude, which forces the ladder of $H$ eigenvalues to terminate at both ends and gives a finite-dimensional unitary representation (with $m=-j,\dots,j$ at fixed $c=j(j+1)$). In contrast, for $\sigma=-1$, the right hand side grows as $m^{2}$, so there is no analogous obstruction: the ladder can extend indefinitely, and any nontrivial unitary representation must be infinite-dimensional.
2.4 CFTs as scaling limits of generalized spin chains
Anyonic chains are generalized spin chains in 1+1 dimensions. A key motivation for studying them is their convergence to minimal model CFTs in their scaling limits Zini and Wang (2018). While numerical simulations provide physical proof of their convergence Feiguin et al. (2007), a mathematically rigorous derivation is still widely open. The scaling limit as defined in Zini and Wang (2018) is a double limit. In the first step, an energy cutoff is prescribed for each system size, and a limit is taken to obtain low energy states and a local observable algebra. Then different system sizes are connected by matching up the low energy states.
Our long-term goal is to develop an analogous scaling limit of the fuzzy sphere Hamiltonians Zhu et al. (2023) reviewed below. Establishing this in a mathematically precise way in the sense of Zini and Wang (2018) would be very challenging, and we leave it to a future publication. Instead, in Sec. 6 we construct a sequence of finite-dimensional non-unitary representations of $so(3,2)$ in terms of the fuzzy sphere density modes. As described above, these representations can also be converted into unitary representations of $so(5)$.
3 Fuzzy sphere review
Fuzzy sphere Hamiltonians exhibit 3d CFT–like spectra with only a handful of spinful fermions Zhu et al. (2023). Their Hamiltonians are naturally written in terms of electron density operators projected to the lowest Landau level and preserve an exact $SO(3)$ symmetry, as we review here.
3.1 Hilbert space and fermionic density fields
The fuzzy sphere model describes spin-$\tfrac{1}{2}$ electrons $\bm{\psi}(\Omega)=(\psi_{\uparrow}(\Omega),\psi_{\downarrow}(\Omega))$ moving on the surface of a sphere with radius $R$ in the presence of a magnetic monopole with flux $4\pi s$ Zhu et al. (2023). The monopole background quantizes the eigenstates of the system into Landau levels. When all interactions are much smaller than the Landau level spacing, we can restrict to the lowest Landau level (LLL). The LLL contains $2s+1$-fold degenerate orbitals for each spin, labeled by $m=-s,\dots,s$. The electron field $\psi_{\sigma}(\Omega)$ projected to the LLL may be expanded as
$\psi_{\sigma}(\Omega)=\frac{1}{R}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\,c_{m,\sigma},\qquad\text{ where }\Omega=(\theta,\phi)\in S^{2},\;\sigma\in\{\uparrow,\downarrow\}.$ (3)
The fermion operators $c_{m,\sigma}$ obey the canonical anticommutation relations
$\{c_{m,\sigma},c^{\dagger}_{m^{\prime},\sigma^{\prime}}\}=\delta_{\sigma,\sigma^{\prime}}\delta_{m,m^{\prime}}$
and $Y^{(s)}_{s,m}(\Omega)$ denotes the monopole harmonics Wu and Yang (1976),
$Y^{(s)}_{s,m}(\Omega)=\sqrt{\frac{(2s+1)!}{(s+m)!(s-m)!}}e^{im\phi}\cos^{s+m}{\left(\frac{\theta}{2}\right)}\sin^{s-m}{\left(\frac{\theta}{2}\right)}.$
[…]
The electron fields projected to the LLL (3) no longer anti-commute at different locations Fardelli et al. (2024):
$\{\psi_{\sigma}(\Omega),\psi^{\dagger}_{\sigma}(\Omega^{\prime})\}=K(\Omega,\Omega^{\prime}),\;\text{ where}$ (4)
$K(\Omega,\Omega^{\prime})=\frac{1}{R}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\left(Y^{(s)}_{s,m}(\Omega^{\prime})\right)^{*}=\frac{2s+1}{R^{2}}e^{2is\alpha(\Omega,\Omega^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\text{ and }$
[…]
The real phase $\alpha(\Omega,\Omega^{\prime})=-\alpha(\Omega^{\prime},\Omega)$ is anti-symmetric as a direct consequence of $K(\Omega^{\prime},\Omega)=K(\Omega,\Omega^{\prime})^{*}$. A derivation of this relation is given in Appendix A.
To relate the sphere radius $R$ to the monopole strength $s$, we use the magnetic field of the monopole. A monopole with charge $s$ produces a uniform radial magnetic field on $S^{2}$ of magnitude $B=s/R^{2}$. Throughout the paper, we fix $R^{2}=2s+1$ so that $B=\frac{s}{2s+1}$ stays finite in the thermodynamic limit $s\to\infty$.
Density operators are defined as bilinears of the fermion field,
$n^{A}(\Omega)=\sum_{\sigma,\lambda\in\{\uparrow,\downarrow\}}\psi^{\dagger}_{\sigma}(\Omega)A_{\sigma\lambda}\,\psi_{\lambda}(\Omega),$
where $A$ is a $2\times 2$ matrix acting on the spin. In particular, we will make frequent use of
$n^{0}(\Omega)=\sum_{\lambda=\uparrow,\downarrow}\psi_{\lambda}^{\dagger}(\Omega)\psi_{\lambda}(\Omega)\qquad\text{and}\qquad n^{\alpha}(\Omega)=\sum_{\lambda,\eta\in\{\uparrow,\downarrow\}}\psi^{\dagger}_{\lambda}(\Omega)\,(\sigma^{\alpha})_{\lambda\eta}\,\psi_{\eta}(\Omega),\qquad\alpha=x,y,z,$
with $\sigma^{\alpha}$ denoting the Pauli matrices. The $n^{A}(\Omega)$ fields can be expanded in spherical harmonics and the corresponding density modes $n^{A}_{l,m}$ defined by
$n^{A}(\Omega)$ $=\frac{1}{2s+1}\sum_{l=0}^{2s}\sum_{m=-l}^{l}n^{A}_{l,m}Y_{l,m}(\Omega),\qquad n^{A}_{l,m}=(2s+1)\int d\Omega\;Y^{*}_{l,m}(\Omega)n^{A}(\Omega).$
After writing $\psi_{\sigma}(\Omega)$ in the orbital basis, these mode operators admit the explicit bilinear form
$n^{A}_{l,m}$ $=(2s+1)\sqrt{2l+1}\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}\sum_{m_{1}=-s}^{s}(-1)^{s+m_{1}}\begin{pmatrix}s&s&l\\ m_{1}&m-m_{1}&-m\end{pmatrix}\sum_{\sigma,\lambda}c^{\dagger}_{m_{1},\sigma}c_{m_{1}-m,\lambda}A_{\sigma\lambda}$ (5)
[…]
3.2 Hamiltonian and symmetries
The fuzzy sphere Hamiltonian that realizes the 3d Ising CFT is given by Zhu et al. (2023); Hofmann et al. (2024)
$H=\int d\Omega\left(U_{0}\left(n^{0}(\Omega)^{2}-n^{z}(\Omega)^{2}\right)+U_{1}\left(n^{0}(\Omega)\nabla^{2}n^{0}(\Omega)-n^{z}(\Omega)\nabla^{2}n^{z}(\Omega)\right)-hn^{x}(\Omega)\right)$ (6)
Expanding in the density modes $n^{A}_{l,m}$, this becomes
$H=-hn^{x}_{0,0}+\sum_{l=0}^{2s}\sum_{m=-l}^{l}(-1)^{m}\left(U_{0}+U_{1}\frac{l(l+1)}{2s+1}\right)\left(n^{0}_{l,m}n^{0}_{l,-m}-n^{z}_{l,m}n^{z}_{l,-m}\right),$ (7)
where we used the hermiticity property
$\left(n^{A}_{l,m}\right)^{\dagger}=(-1)^{m}n^{A^{\dagger}}_{l,-m}.$
The filling (total number of electrons) is
$N_{\text{el}}=\sum_{m=-s}^{s}(n^{\uparrow}_{m}+n^{\downarrow}_{m})=(-1)^{2s}n^{0}_{0,0},$
and commutes with every $n^{A}_{l,m}$. To study the 3d Ising CFT in the fuzzy sphere model, one can restrict to the sector with filling $N_{\text{el}}=N$. For large $h$, the model is in the trivial phase with a unique ground state. For small $h$, it enters a quantum Hall ferromagnet phase with two degenerate ground states, spontaneously breaking the $\mathbb{Z}_{2}$ spin-flip symmetry acting as $c_{m,\uparrow}\leftrightarrow c_{m,\downarrow}$. In the $(h,U_{0})$ phase diagram (with $U_{1}$ fixed), these phases are separated by a critical line along which the low-energy spectrum organizes into conformal multiplets consistent with the 3d Ising CFT. Because finite-size effects vary appreciably along this line, it is useful in practice to tune both $h$ and $U_{1}$.
Crucially, the Hamiltonian preserves an exact $SO(3)$ rotational symmetry, which appears to be important for extracting CFT data already at small system sizes. The $SO(3)$ generators act on the orbital indices in the spin-$s$ representation and are proportional to the $n^{0}_{1m}$ density modes:
$J_{z}$ $=(s+1)\frac{(-1)^{2s}}{\sqrt{3}}n^{0}_{1,0}=\sum_{m=-s}^{s}m\left(c^{\dagger}_{m,\uparrow}c_{m,\uparrow}+c^{\dagger}_{m,\downarrow}c_{m,\downarrow}\right),$ (8)
$J_{\pm}$ $=\pm(s+1)(-1)^{2s}\sqrt{\frac{2}{3}}n^{0}_{1,\pm 1}=\sum_{m=-s}^{s}\sqrt{(s\mp m)(s\pm m+1)}\left(c^{\dagger}_{m\pm 1,\uparrow}c_{m,\uparrow}+c^{\dagger}_{m\pm 1,\downarrow}c_{m,\downarrow}\right).$
Using the commutation relations (9) reviewed below, it is straightforward to check that $H,J_{z}]=[H,J_{\pm}]=0$. Since spatial rotations only act on the orbital sector $V^{s}\cong\mathbb{C}^{2s+1}$ (not on the internal spin/flavor indices), the relevant representation theory is that of operators on $V^{s}$, i.e. $\mathrm{End}(V_{s})$. Under the adjoint $so(3)$ action, $\mathrm{End}(V_{s})$ decomposes as
$\mathrm{End}(V_{s})\cong V^{s}\otimes\overline{V^{s}}=\oplus_{l=0}^{2s}\;V^{l}.$
Accordingly, any rotation-invariant fuzzy sphere Hamiltonian has eigenstates that organize into $SO(3)$ multiplets labeled by $l$.
Beyond the Ising CFT, similar fuzzy sphere Hamiltonians have been used to realize a variety of 3d CFTs, including the free real scalar CFT He (2025); Taylor et al. (2025), the $O(N)$ and Wilson-Fisher CFTs Dey et al. (2025); Guo et al. (2025), DQCPs with emergent $SO(5)$ symmetry Zhou et al. (2024); Chen et al. (2024), the non-unitary Yang-Lee CFT Fan et al. (2025); Elias Miró and Delouche (2025); Arguello Cruz et al. (2025), a superconformal Ising CFT Tang et al. (2025), and the free fermion Zhou et al. (2025a) and Chern-Simons-matter CFTs Zhou and He (2025); Zhou et al. (2025b).
3.3 Conformal generators for the fuzzy sphere Hamiltonian
Refs. Fardelli et al. (2024) and Fan (2024) found a lattice precursor of the conformal generators $\Lambda_{m}=K_{m}+P_{m}$ with $m\in\{0,\pm 1\}$ ($\Lambda_{0}=\Lambda_{z}$), where $K_{m}$ and $P_{m}$ generate translations and special conformal transformations as in (1). The lattice operator corresponding to $\Lambda_{m}$ is obtained by taking the $l=1$ spherical-harmonic moment of the Hamiltonian density $H(\Omega)$ given in (6):
$\Lambda_{m}=\int d\Omega\;Y_{1,m}(\Omega)\,H(\Omega),$
All other conformal generators follow from this identification. In particular, $P_{z}=H+[H,\Lambda_{z}]$ and $K_{z}=H-[H,\Lambda_{z}]$. Because $H(\Omega)$ is invariant under $\mathbb{Z}_{2}$ spin-flip, these lattice precursors inherit the same symmetry. In practice, their overlap with the CFT generators can be improved by tuning couplings ($h$, $U_{0}$, $U_{1}$) to suppress descendant contributions Fardelli et al. (2024, 2026).
At finite $s$, this identification does not satisfy the $so(3,2)$ algebra exactly. For example, the commutator $P_{z},K_{z}]=2D$ (equivalently, $\Lambda_{z},[D,\Lambda_{z}]]=4D$) holds only within a low-energy subspace, as demonstrated numerically in Fardelli et al. (2024). The energy window over which this relation is well satisfied grows with $\sqrt{s}$, consistent with conformal symmetry emerging in the continuum limit.
4 Fuzzy sphere algebra
The angular momentum modes of the electron density obey an interesting algebra that can be viewed as a spin-enriched version of the Girvin-MacDonald-Platzman (GMP) algebra on the sphere He (2025). Section 4.1 reviews this algebra and collects results that use its concrete fermion-bilinear realization; Section 4.2 treats the commutator relations as defining an abstract bracket and proves representation-independent properties.
4.1 Commutation relations of the density modes
The mode operators (5) satisfy the following commutation relations He (2025):
$\left[n^{A}_{l,m},n^{B}_{l^{\prime},m^{\prime}}\right]=$ $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ even}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{[A,B]}_{L,m+m^{\prime}}$ (9)
- $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ odd}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{\{A,B\}}_{L,m+m^{\prime}}$,
where
$h_{L}(s,l,l^{\prime})=(2s+1)\sqrt{(2l+1)(2l^{\prime}+1)(2L+1)}\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix}\frac{\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}\begin{pmatrix}s&s&l^{\prime}\\ s&-s&0\end{pmatrix}}{\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}}$. (10)
We derive the density–mode commutator in Appendix B by rewriting the mode operators in a basis of fuzzy spherical harmonics $\hat{Y}_{l,m}$ Madore (1992); Grosse et al. (1996); Iso et al. (2001). The matrices $\hat{Y}_{l,m}$ furnish a $SO(3)$ covariant basis of the matrix algebra $\text{Mat}_{2s+1}$. Their matrix entries are given by
$\left(\hat{Y}_{l,m}\right)_{ab}=(-1)^{s-a}\sqrt{2l+1}\begin{pmatrix}s&s&l\\ a&-b&-m\end{pmatrix}$,
and they satisfy
[…]
Up to a prefactor, the density modes can be written as fermion bilinears with orbital matrix $\hat{Y}_{l,m}$:
$n^{A}_{l,m}=\mathcal{N}_{s,l}\sum_{m_{1},m_{2}=-s}^{s}\left(\hat{Y}_{l,m}\right)_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}A_{\sigma\lambda}c_{m_{2},\lambda},\qquad\text{with }\mathcal{N}_{s,l}=(2s+1)\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}$. (11)
We suppress explicit summation over the spin indices in what follows.
The fermion bilinears $E_{ij}:=c^{\dagger}_{i}c_{j}$ with combined orbital-spin indices $i=(m_{1},\sigma)$, $j=(m_{2},\lambda)$ generate $u(2N)$ ($N=2s+1$) with the standard commutation relations
$\left[E_{ij},E_{kl}\right]=\delta_{jk}E_{il}-\delta_{li}E_{kj}.$
In their fermionic representation (11), each mode $n^{A}_{l,m}$ is a linear combination of $E_{ij}$. Using the completeness of the orbital and spin bases, this change of basis is invertible; thus, the two sets span the same vector space and generate the same matrix Lie algebra $u(2N)$.
Besides the commutation relations, there are additional polynomial relations among the fermionic density operators. Using the LLL anticommutation relation, the squared density field can be written as He (2025)
$n^{A}(\Omega)^{2}$ $=\sum_{\alpha\beta\gamma\delta}\psi^{\dagger}_{\alpha}(\Omega)A_{\alpha\beta}\psi_{\beta}(\Omega)\psi^{\dagger}_{\gamma}(\Omega)A_{\gamma\delta}\psi_{\delta}(\Omega)$ (12)
$=\sum_{\alpha\beta\gamma\delta}A_{\alpha\beta}A_{\gamma\delta}\left(\psi^{\dagger}_{\alpha}(\Omega)K(\Omega,\Omega)\delta_{\beta\gamma}\psi_{\delta}(\Omega)-\psi^{\dagger}_{\alpha}(\Omega)\psi^{\dagger}_{\gamma}(\Omega)\psi_{\beta}(\Omega)\psi_{\delta}(\Omega)\right)$
$=n^{A^{2}}(\Omega)-\sum_{\alpha\beta\gamma\delta}A_{\alpha\beta}A_{\gamma\delta}\psi^{\dagger}_{\alpha}(\Omega)\psi^{\dagger}_{\gamma}(\Omega)\psi_{\beta}(\Omega)\psi_{\delta}(\Omega)$
For $A\in\{\sigma^{x},\sigma^{y},\sigma^{z}\}$ we have $A^{2}=\mathbb{I}$, and the quartic term can be eliminated in favor of $n^{0}(\Omega)^{2}$ using
$n^{0}(\Omega)^{2}=n^{0}(\Omega)-\sum_{\alpha,\beta}\psi^{\dagger}_{\alpha}(\Omega)\psi^{\dagger}_{\beta}(\Omega)\psi_{\alpha}(\Omega)\psi_{\beta}(\Omega),$
which yields
$n^{x}(\Omega)^{2}=n^{y}(\Omega)^{2}=n^{z}(\Omega)^{2}=2n^{0}(\Omega)-n^{0}(\Omega)^{2}.$
Integrating over $\Omega$ gives the corresponding quadratic relations among density modes:
$\sum_{l,m}(-1)^{m}n^{\alpha}_{l,m}\,n^{\alpha}_{l,-m}=2n^{0}_{0,0}-\sum_{l,m}(-1)^{m}n^{0}_{l,m}\,n^{0}_{l,-m},\hskip 18.49988pt\alpha=x,y,z.$ (13)
4.2 Jacobi identity and commuting subalgebras
Does the mode algebra, viewed abstractly through the commutation relations (9), define a genuine Lie algebra? Let $\mathcal{A}_{s}=\{\tilde{n}^{A}_{l,m}\}_{l=0,|m|\leq l}$ be the algebra spanned by formal generators $\tilde{n}^{A}_{l,m}$, equipped with the bracket defined by (9). A priori this bracket need not satisfy Jacobi. To verify Jacobi, we use the fermionic realization. Let $\rho:\mathcal{A}_{s}\to\operatorname{End}(\mathcal{H}_{\rm Fock})$ be the map that sends the abstract generators to the corresponding lowest-Landau-level bilinears,
$\rho(\tilde{n}^{A}_{l,m}):=n^{A}_{l,m}\hskip 18.49988ptA\in\{\mathbb{I},\sigma^{x},\sigma^{y},\sigma^{z}\},$
where $n^{A}_{l,m}$ on the right-hand side is the explicit operator defined in (5). Since $\operatorname{End}(\mathcal{H}_{\rm Fock})$ is an associative algebra, its commutator bracket satisfies Jacobi. Hence for all $x,y,z\in\mathcal{A}_{s}$ the Jacobiator
$J(x,y,z):=[x,[y,z]]+[y,[z,x]]+[z,[x,y]]$
obeys $\rho(J(x,y,z))=0$. Therefore, if $\rho$ is faithful, it follows that $J(x,y,z)=0$ already in $\mathcal{A}_{s}$, i.e. the commutator (9) defines a genuine Lie algebra.
We prove faithfulness by showing that the representation matrices $\{n^{A}_{l,m}\}$ are linearly independent. If the linear relation $\sum_{i}\alpha_{i}n^{A_{i}}_{l_{i},m_{i}}=0$ holds, then $\sum_{i}\alpha_{i}[n^{A_{i}}_{l_{i},m_{i}},c^{\dagger}_{r,\eta}]=0$ also holds. Inserting the fermionic bilinear form (11) into the latter equation and using
$[c^{\dagger}_{m_{1},\sigma}c_{m_{2},\lambda},c^{\dagger}_{r,\eta}]=\delta_{r,m_{2}}\delta_{\lambda,\eta}c^{\dagger}_{m_{1},\sigma}$
results in
$\sum_{i}\alpha_{i}\left[n^{A_{i}}_{l_{i},m_{i}}),c^{\dagger}_{r,\eta}\right]=\sum_{i}\alpha_{i}\,\mathcal{N}_{s,l_{i}}\sum_{m_{1}}\left(\hat{Y}_{l_{i},m_{i}}\right)_{m_{1},r}A^{i}_{\sigma_{i},\eta}c^{\dagger}_{m_{1},\sigma_{i}}=0.$
The creation operators $c^{\dagger}_{m_{1},\sigma_{i}}$ are linearly independent, requiring the expression above to vanish for each $m_{1},r,\sigma,\eta$. Because the $A_{i}$ form a basis of Mat2 and the fuzzy spherical harmonics $\hat{Y}_{l,m}$ a basis of Mat2s+1, this implies $\alpha_{i}=0$ for all $i$. Therefore, the $\{n^{A}_{l,m}\}$ are linearly independent and $\rho$ is faithful.
It would be very interesting to characterize this mode algebra more intrinsically, i.e. to find a presentation of each fuzzy sphere algebra $\mathcal{A}_{s}$ with generators $\{\tilde{n}^{A}_{l,m}\}$ and a minimal set of relations that define the algebra $\mathcal{A}_{s}$, potentially including identities between squared modes as in (13). If one can endow $\mathcal{A}_{s}$ with an associative multiplication $x\cdot y$ such that, for the generators, the induced commutator $x,y]=x\cdot y-y\cdot x$ agrees with the bracket in (9), then the Jacobi identity would follow automatically: the commutator in any associative algebra is a Lie bracket.
The commutator (9) comes with strong angular-momentum selection rules, and these imply the existence of abelian subalgebras at the level of the abstract algebra. In the spinless sector spanned by $\{\tilde{n}^{0}_{l,m}\}$, one finds three families of mutually commuting modes Dora and Balram (2025):
$\left[\tilde{n}^{0}_{l,0},\tilde{n}^{0}_{l^{\prime},0}\right]=0,\qquad\left[\tilde{n}^{0}_{l,l},\tilde{n}^{0}_{l^{\prime},l^{\prime}}\right]=0,\qquad\left[\tilde{n}^{0}_{l,-l},\tilde{n}^{0}_{l^{\prime},-l^{\prime}}\right]=0.$
The first commutator vanishes because the 3j symbol $\begin{pmatrix}l&l^{\prime}&L\\ 0&0&0\end{pmatrix}$ that appears in (9) vanishes for $l+l^{\prime}+L$ odd. The second commutator can only contain modes with $m=l+l^{\prime}\leq L$, but the selection rules of the 3j symbol require $L\leq l+l^{\prime}$, so the only possibility is $L=l+l^{\prime}$. This choice makes $l+l^{\prime}+L$ even, and so the term is excluded by the $l+l^{\prime}+L$ odd constraint in the sum. A similar argument shows that the third commutator vanishes. Therefore, the rank of the subalgebra generated by the $\tilde{n}^{0}_{l,m}$ is $N=2s+1$. For the full algebra generated by the $\tilde{n}^{A}_{l,m}$, the rank doubles to $2N$, and a natural maximal commuting subalgebra is generated by $\{\tilde{n}^{0}_{l,0},\,\tilde{n}^{z}_{l,0}\}$.
Finally, a separate question is whether this algebra admits a nontrivial central extension; we briefly comment on this in Appendix C.
5 Thermodynamic limit of the density mode algebra
The fuzzy sphere geometry has two fundamentally different large-$s$ limits Chu et al. (2001); Steinacker (2024): On the one hand, if we zoom into a microscopic patch near the north pole, the algebra locally approaches the fuzzy plane: modes with angular momentum $l\sim\sqrt{s}$ have finite planar momentum $q\sim l/R$ and become generators of the GMP algebra. On the other hand, for small angular momenta $l\ll\sqrt{s}$, the non-commutativity vanishes as $s\to\infty$ and the fuzzy sphere becomes a commutative sphere. For the critical fuzzy-sphere Hamiltonians, we expect the low-energy sector to be controlled by this latter, semiclassical regime.
5.1 Quantum planar limit of large-$l$ density modes
Following Appendix E of Matsuura and Sugino (2016), we approximate the fuzzy spherical harmonics matrices $\hat{Y}_{l,m}$ that appear in the density modes (11), focusing on the regime $l\sim\sqrt{s}$. In the large-$s$ limit with $l\ll s$, we may use the asymptotic expansion of Clebsch-Gordan coefficients (Varshalovich et al., 1988, §8.9.1)
$\left(\hat{Y}_{l,m}\right)_{m_{1},m_{2}}=(-1)^{s-m_{2}}\,C^{lm}_{sm_{1},s(-m_{2})}\ \approx\ \delta_{m,m_{1}-m_{2}}\,\sqrt{\frac{4\pi}{2s+1}}\,Y_{lm}(\theta,0),$ (14)
where the small polar angle $\theta$ is related to the row index $m_{1}$ by
$\sin\frac{\theta}{2}=\sqrt{\frac{m_{1}+\tfrac{1}{2}}{2s+1}}.$ (15)
(Here the north-pole patch corresponds to $m_{1}\ll s$, so $\theta\ll 1$.) When $l\sim\sqrt{s}$, one can further use the Bessel function approximation of spherical harmonics in the limit
$l\to\infty,\qquad m\to\infty,\qquad\theta\to 0,\qquad\text{with }l\theta\ \text{fixed},$
which yields (Varshalovich et al., 1988, §5.12.3)
$Y_{lm}(\theta,0)\ \approx\ \sqrt{\frac{l}{2\pi}}\,(-1)^{m}\,J_{m}(l\theta).$ (16)
The resulting approximation for the fuzzy spherical harmonics is
$\left(\hat{Y}_{l,m}\right)_{m_{1},m_{2}}\ \approx\delta_{m,m_{1}-m_{2}}\,\sqrt{\dfrac{2l}{2s+1}}\,(-1)^{m}\,J_{m}(l\theta).$ (17)
Next, it is useful to rewrite (17) in a form that makes the connection to the noncommutative plane clearer Matsuura and Sugino (2016),
$\left(\hat{Y}_{l,m}\right)_{m_{1},m_{2}}\ \approx\ (-i)^{m}\,\sqrt{\frac{2l}{2s+1}}\int_{0}^{2\pi}\frac{d\phi}{2\pi}\,e^{im\phi}\left(e^{\,i\mathbf{q}(\phi)\cdot\hat{\mathbf{x}}}\right)_{m_{1},m_{2}},\hskip 18.49988pt|\mathbf{q}(\phi)|=\frac{l}{R},$ (18)
where $e^{\,i\mathbf{q}\cdot\hat{\mathbf{x}}}$ is the plane-wave operator on the quantum plane, with noncommuting coordinates
$\hat{x}_{1},\hat{x}_{2}]=i\Theta,\hskip 18.49988pt\Theta\sim\frac{R^{2}}{2s+1}\sim\frac{1}{2B}.$ (19)
Inserting (18) into the definition of the density modes (11) gives
$n^{A}_{l,m}\ \approx\ \kappa_{s,l}\int_{0}^{2\pi}\frac{d\phi}{2\pi}\,e^{im\phi}\,\rho^{A}_{\mathbf{q}(\phi)},\hskip 18.49988pt|\mathbf{q}(\phi)|=\frac{l}{R},$ (20)
with some prefactor $\kappa_{s,l}$, and where the planar density modes are
$\rho^{A}_{\mathbf{q}}:=\sum_{m_{1},m_{2}=0}^{\infty}\left(e^{\,i\mathbf{q}\cdot\hat{\mathbf{x}}}\right)_{m_{1},m_{2}}\,c^{\dagger}_{m_{1},\sigma}\,A_{\sigma\lambda}\,c_{m_{2},\lambda}.$ (21)
Thus the large-$l$ spherical density modes become angular Fourier components of planar momentum modes $\rho^{A}_{\mathbf{q}}$ with $|\mathbf{q}|=l/R$.
The operators $e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}$ satisfy the algebra following from (19),
$e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}\,e^{i\mathbf{p}\cdot\hat{\mathbf{x}}}=\exp\!\Big(-\frac{i}{2}\Theta\,(\mathbf{q}\times\mathbf{p})\Big)\,e^{i(\mathbf{q}+\mathbf{p})\cdot\hat{\mathbf{x}}},\hskip 18.49988pt\mathbf{q}\times\mathbf{p}:=q_{1}p_{2}-q_{2}p_{1}.$ (22)
Using the fermion bilinear expansion (21), this implies the spin-enriched planar GMP algebra
$\left[\rho^{A}_{\mathbf{q}},\rho^{B}_{\mathbf{p}}\right]=\cos\!\Big(\frac{\Theta}{2}\,(\mathbf{q}\times\mathbf{p})\Big)\,\rho^{[A,B]}_{\mathbf{q}+\mathbf{p}}+i\sin\!\Big(\frac{\Theta}{2}\,(\mathbf{q}\times\mathbf{p})\Big)\,\rho^{\{A,B\}}_{\mathbf{q}+\mathbf{p}}$ (23)
The spinless subalgebra ($A=B=\mathbb{I}$) reduces to the planar GMP algebra Girvin et al. (1985).
From (23), the commutator of the $l\sim\sqrt{s}$ modes in the large-$s$ limit can be written as an angular average of the planar GMP commutator:
[…]
where
$\mathbf{q}(\phi)=\frac{l}{R}(\cos\phi,\sin\phi),\hskip 18.49988pt\mathbf{q}^{\prime}(\phi^{\prime})=\frac{l^{\prime}}{R}(\cos\phi^{\prime},\sin\phi^{\prime}),\hskip 18.49988pt\beta=\frac{\Theta}{2}\big(\mathbf{q}(\phi)\times\mathbf{q}^{\prime}(\phi^{\prime})\big).$
It is convenient to change variables to $\Phi=\frac{\phi+\phi^{\prime}}{2}$ and $\Delta=\phi^{\prime}-\phi$ so that
$e^{im\phi}e^{im^{\prime}\phi^{\prime}}=e^{i(m+m^{\prime})\Phi}\,e^{i(m^{\prime}-m)\Delta/2}.$
In these variables, the scalars $\beta$ and the magnitude $|\mathbf{Q}|:=|\mathbf{q}(\phi)+\mathbf{q}^{\prime}(\phi^{\prime})|$ depend only on $\Delta$ (through $\sin\Delta$ and $\cos\Delta$), whereas the direction of $\mathbf{Q}:=\mathbf{q}(\phi)+\mathbf{q}^{\prime}(\phi^{\prime})$ is
[…]
Writing the inverse angular decomposition (the inverse of (20))
$\rho^{A}_{\mathbf{Q}}\;=\;\sum_{M\in\mathbb{Z}}e^{-iM\varphi_{\mathbf{Q}}}\;n^{A}_{L,M},\hskip 18.49988ptL:=|\mathbf{Q}|\,R,$ (25)
and inserting (25) into (24), the integral over $\Phi$ enforces angular-momentum conservation:
$\int_{0}^{2\pi}\frac{d\Phi}{2\pi}\,e^{i(m+m^{\prime}-M)\Phi}=\delta_{M,m+m^{\prime}}.$
As a result,
[…]
where $L(\Delta)=|\mathbf{q}+\mathbf{q}^{\prime}|\,R=\sqrt{l^{2}+l^{\prime 2}+2ll^{\prime}\cos\Delta}$ and $\beta(\Delta)=\frac{\Theta}{2}\frac{ll^{\prime}}{R^{2}}\sin\Delta$.
We can use the large-$l$ correspondence (20) to rewrite the fuzzy-sphere Hamiltonian (7) in the planar limit. Substituting $n^{A}_{l,m}\mapsto\rho^{A}_{\mathbf{q}}$ (with $|\mathbf{q}|=l/R$) yields
$H\approx\rho^{x}_{\mathbf{0}}+\int d^{2}q\;V(q)\Big(\rho^{z}_{\mathbf{q}}\rho^{z}_{-\mathbf{q}}+\rho^{0}_{\mathbf{q}}\rho^{0}_{-\mathbf{q}}\Big),\hskip 18.49988ptV(q)=U_{0}+U_{1}q^{2}.$ (26)
The first term acts as a uniform field, and the remaining terms describe planar density–density interactions.
5.2 Semiclassical limit of small-$l$ density modes
The low-energy sector of the critical fuzzy-sphere Hamiltonian is dominated by modes with angular momentum $l\ll\sqrt{s}$. In the thermodynamic limit this sector is expected to probe the semiclassical (commutative) limit of the fuzzy sphere, rather than the planar GMP regime discussed above Zhou (2025); Fardelli et al. (2024). The onset of the planar regime at $l\sim\sqrt{s}$ has also been observed numerically Fardelli et al. (2024); He (2025).
In the semiclassical regime $l\ll\sqrt{s}$, the fuzzy spherical harmonics matrices $\hat{Y}_{l,m}$ can be identified with the ordinary spherical harmonics $Y_{l,m}(\Omega)$ on $S^{2}$ Steinacker (2024). Moreover, as $s\to\infty$ the matrix product approaches the pointwise product of functions, while the commutator reduces to the Poisson bracket on the sphere. Therefore,
$\big[n^{A}[f],n^{B}[g]\big]\;=\;n^{[A,B]}[fg]\;+\;\mathcal{O}\!\left(\frac{1}{s}\right)\,i\,n^{\{A,B\}}\!\left[\{f,g\}_{\rm PB}\right],$ (27)
where $n^{A}[f]$ is defined as
$n^{A}[f]=\int d\Omega\;f(\Omega)\psi^{\dagger}_{\sigma}(\Omega)A_{\sigma\lambda}\psi_{\lambda}(\Omega).$
Taking $f=Y_{l,m}$ and $g=Y_{l^{\prime},m^{\prime}}$ with $l,l^{\prime}\ll\sqrt{s}$ yields the semiclassical commutator of the corresponding small-$l$ density modes. Using the standard product decomposition of spherical harmonics, the leading term is
$[n^{A}_{l,m},n^{B}_{l^{\prime},m^{\prime}}]\;\approx\;n^{[A,B]}\!\left[\,Y_{l,m}Y_{l^{\prime},m^{\prime}}\,\right]$ (28)
$=\sum_{L}\sqrt{(2l+1)(2l^{\prime}+1)(2L+1)}\;\begin{pmatrix}l&l^{\prime}&L\\ 0&0&0\end{pmatrix}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-(m+m^{\prime})\end{pmatrix}(-1)^{m+m^{\prime}}\,n^{[A,B]}_{L,\,m+m^{\prime}}$.
Here the first $3j$ symbol enforces the triangle inequalities (in particular $L\leq l+l^{\prime}$, so the result remains in the small-$l$ sector) and is nonzero only when $l+l^{\prime}+L$ is even.
However, this naive semiclassical bracket does not directly reproduce several relations that appear to emerge in the low-energy sector of critical fuzzy-sphere Hamiltonians. For instance, Ref. He (2025) found that, upon restricting to states near the paramagnetic ground state, the operators $n^{+}_{l,m}$ and $n^{-}_{l,m}$ approximately behave as oscillator modes, with $n^{+}_{l,m},n^{-}_{l^{\prime},m^{\prime}}]\approx\delta_{l,l^{\prime}}\delta_{m+m^{\prime},0}n^{z}_{0,0}$. By contrast, the semiclassical commutator (28) generically yields a sum over many $L$ rather than collapsing to a single $L=0$ term. Similarly, the semiclassical commutator (27) does not by itself imply the conformal commutation relations observed numerically in the same low-energy subspace, e.g. $\Lambda_{z},[\Lambda_{z},D]]=4D$ in Fardelli et al. (2024), from the identifications $D\sim H$ and $\Lambda_{z}\sim\int d\Omega\,Y_{1,0}(\Omega)H(\Omega)$. Capturing these emergent relations likely requires a more refined scaling limit, as discussed in Sec. 2.4.
6 Hopf algebra construction of $so(3,2)$ representations
Establishing analytically how conformal symmetry emerges in the low-energy continuum limit of the fuzzy sphere Hamiltonian appears too difficult. Instead, we ask a much simpler question: does there exist an $so(3,2)$ representation whose generators are density modes $n^{A}_{l,m}$? We find that such a representation exists in the minimal nontrivial case of $s=\tfrac{1}{2}$, but does not persist for $s>\tfrac{1}{2}$. One can nevertheless extend the $s=\tfrac{1}{2}$ representation to larger systems via a Hopf algebra coproduct. However, this coproduct maps a spin-$s$ irrep of $so(3)$ to a tensor product of spin-$k$ and spin-$l$ irreps, so it splits a single fuzzy sphere into two fuzzy spheres. Therefore, it does not capture the thermodynamic limit relevant for our critical Hamiltonians, where one increases $s$ to obtain a single larger fuzzy sphere.
6.1 $so(3,2)$ algebra representation of $s=\tfrac{1}{2}$ density modes
We identify the rotation generators of the $so(3,2)$ algebra as usual with
$J_{z}$ $=\frac{-2}{\sqrt{3}}n^{0}_{1,0},\hskip 18.49988ptJ_{\pm}=\mp 2\sqrt{\frac{2}{3}}n^{0}_{1,\pm 1}.$
Then we propose the following identifications for the $so(2,1)$ generators (as defined in Section 3.3):
$D_{z}=-3n^{z}_{0,0},\qquad D_{+}=i\sqrt{3}n^{+}_{1,0},\qquad D_{-}=i\sqrt{3}n^{-}_{1,0}.$
Note that the $so(2,1)$ algebra acts on the spin degrees of freedom, corresponding to $\sigma^{z}$ and $i\sigma^{\pm}$ operators. The remaining four ladder operators follow as
$R_{\pm}=\pm[J_{\pm},D_{\pm}]=\mp i\sqrt{6}n^{\pm}_{1,\pm 1},\qquad S_{\pm}=\pm[J_{\mp},D_{\pm}]=\mp i\sqrt{6}n^{\pm}_{1,\mp 1}.$
It is straightforward to check that the defining $so(3,2)$ commutation relations (2) are satisfied.
[…]
To obtain a unitary $so(5)$ representation instead of the non-unitary $so(3,2)$ one, it is enough to remove the factors of $i$ in $D_{\pm}$. With
$D_{z}=-n^{z}_{0,0},\qquad D_{+}^{so(5)}=\sqrt{3}n^{+}_{1,0},\qquad D_{-}^{so(5)}=\big(D_{-}^{so(5)}\big)^{\dagger}=\sqrt{3}n^{-}_{1,0},$
the generators $D_{z}$ and $D_{\pm}$ now close into an $so(3)$ subalgebra. The $R^{so(5)}_{\pm}$ and $S^{so(5)}_{\pm}$ operators are then defined by the same commutation relations as before, but with $D_{\pm}$ replaced with $D_{\pm}^{so(5)}$.
In Appendix D, we show that this so(3,2)so(3,2)/$so(5)$ representation built from the $n^{A}_{l,m}$ density modes at $s=\tfrac{1}{2}$ does not extend to larger $s$, even if we allow an arbitrary enlargement of the internal spin degrees of freedom. For example, it is easy to see that the commutator $D_{+},D_{-}]$ produces an additional $n^{z}_{2,0}$ term when $s>\tfrac{1}{2}$.
6.2 Larger $so(3,2)$ representation from $so(3)$ equivariant coproduct
[…]
where $C^{Ji}_{Kk,Ll}$ are Clebsch-Gordon coefficients. The coproduct is nonzero only when $\rho^{J}$ appears in the tensor product $\rho^{K}\otimes\rho^{L}$ (equivalently $|K-L|\leq J\leq K+L$). It is a homomorphism, i.e. $\Delta_{KL}(M_{1}M_{2})=\Delta_{KL}(M_{1})\Delta_{KL}(M_{2})$.
The coproduct is $so(3)$-equivariant with respect to the left action: for any generator $X\in so(3)$ and any $M\in\text{Mat}(2J{+}1)$,
$\Delta_{KL}\!\big(\rho^{J}(X)\,M\big)=\Big(\rho^{K}(X)\otimes\mathbb{I}+\mathbb{I}\otimes\rho^{L}(X)\Big)\,\Delta_{KL}(M),$
and analogously for the right action. Equivariance implies that the coproduct of an $so(3)$ generator itself is
$\Delta_{K,L}\!\left(\rho^{J}(X)\right)$ $=P_{J}\Big(\rho^{K}(X)\otimes\mathbb{I}+\mathbb{I}\otimes\rho^{L}(X)\Big)P_{J},\qquad\text{where }P_{J}=\Delta_{K,L}(\mathbb{I})$
is the projector onto the spin-$J$ irrep inside $K\otimes L$. The equivariant coproduct of the rotation generators is thus the usual Lie coproduct, followed by a projection onto the spin $J$ channel. One can also define a counit and antipode, leading to a Hopf algebra Balachandran and Kurkcuoglu (2004) (strictly speaking, only when $J=K=L$).
To apply $\Delta_{KL}$ to the density modes $n^{A}_{l,m}$, we write them in their fermionic representation (11) and apply the coproduct to the orbital matrix. For our purposes, we choose $J=\tfrac{1}{2}$, $K=s$, and $L=s+\tfrac{1}{2}$, so that $\Delta_{s,s+\frac{1}{2}}$ maps an operator on the spin-$1/2$ fuzzy sphere to an operator on a spin-$(s+\tfrac{1}{2})$ fuzzy sphere coupled to a spin-$s$ sphere:
$\Delta_{s,s+1/2}(n^{A}_{l,m})=\sum_{m_{1},m_{2}=-s}^{s}\;\sum_{n_{1},n_{2}=-s-1/2}^{s+1/2}c^{\dagger}_{m_{1}n_{1},\sigma}\,\big(\Delta_{s,s+1/2}(\hat{Y}_{l,m})\big)_{m_{1}n_{1},m_{2}n_{2}}\,c_{m_{2}n_{2},\lambda}A_{\sigma,\lambda}.$
While this coproduct gives us a $so(3,2)$ representation defined on a larger Hilbert space, it is different from the thermodynamic limit of the critical fuzzy sphere models, which takes $s\to\infty$ within a single $V^{s}$. Using the standard coproduct of Lie algebras leads to a similar mismatch, see Appendix E. Nevertheless, it seems interesting that this coproduct induces a nontrivial $so(3,2)$ representation on a tensor product Hilbert space.
Acknowledgments
We thank Yin-chen He for his insightful comments on an early draft. Z.W. is partially supported by ARO MURI contract W911NF-20-1-0082. L.E. is supported by the Walter Burke Institute for Theoretical Physics at Caltech. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendices
Appendix A Derivation of the fermionic field anti-commutator
Here we derive the anti-commutator of the fermionic fields (3). By definition of the fermionic fields (3), we have
$\{\psi_{\sigma}(\theta,\phi),\psi^{\dagger}_{\sigma}(\theta,\phi^{\prime})\}=\frac{1}{2s+1}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\theta,\phi)\left(Y^{(s)}_{s,m}(\theta^{\prime},\phi^{\prime})\right)^{*}$
$=\sum_{m=-s}^{s}\frac{(2s)!}{(s+m)!(s-m)!}e^{im(\phi-\phi^{\prime})}\left(\cos\left(\frac{\theta}{2}\right)\cos\left(\frac{\theta^{\prime}}{2}\right)\right)^{s+m}\left(\sin\left(\frac{\theta}{2}\right)\sin\left(\frac{\theta^{\prime}}{2}\right)\right)^{s-m}$
$=e^{-is(\phi-\phi^{\prime})}\sum_{k=0}^{2s}\frac{(2s)!}{k!(2s-k)!}\left(e^{i(\phi-\phi^{\prime})}\cos\left(\frac{\theta}{2}\right)\cos\left(\frac{\theta^{\prime}}{2}\right)\right)^{k}\left(\sin\left(\frac{\theta}{2}\right)\sin\left(\frac{\theta^{\prime}}{2}\right)\right)^{2s-k}$
$=e^{-is(\phi-\phi^{\prime})}\sum_{k=0}^{2s}\binom{2s}{k}\left(e^{i(\phi-\phi^{\prime})}\cos\left(\frac{\theta}{2}\right)\cos\left(\frac{\theta^{\prime}}{2}\right)\right)^{k}\left(\sin\left(\frac{\theta}{2}\right)\sin\left(\frac{\theta^{\prime}}{2}\right)\right)^{2s-k}$
$=e^{-is(\phi-\phi^{\prime})}\left(e^{i(\phi-\phi^{\prime})}\cos\left(\frac{\theta}{2}\right)\cos\left(\frac{\theta^{\prime}}{2}\right)+\sin\left(\frac{\theta}{2}\right)\sin\left(\frac{\theta^{\prime}}{2}\right)\right)^{2s}$
[…]
The phase $e^{i\chi(\theta,\phi,\theta^{\prime},\phi^{\prime})}$ of this term can be derived as
[…]
where the numerator is the imaginary part of the term and the denominator is the real part. So overall, we have
$\{\psi_{\sigma}(\theta,\phi),\psi^{\dagger}_{\sigma}(\theta,\phi^{\prime})\}=e^{2is\alpha(\theta,\phi,\theta^{\prime},\phi^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\qquad\alpha(\theta,\phi,\theta^{\prime},\phi^{\prime})=-\frac{\phi-\phi^{\prime}}{2}+\chi(\theta,\phi,\theta^{\prime},\phi^{\prime}).$
The real phase $\alpha$ is anti-symmetric under exchanging $\theta\leftrightarrow\theta^{\prime}$ and $\phi\leftrightarrow\phi^{\prime}$.
Appendix B Derivation of the density mode commutator using fuzzy spherical harmonics
To derive the density mode commutator (9), it is helpful to rewrite the density modes as in (11),
$n^{A}_{l,m}=\mathcal{N}_{s,l}\sum_{m_{1},m_{2}=-s}^{s}\left(\hat{Y}_{l,m}\right)_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}A_{\sigma\lambda}c_{m_{2},\lambda},\qquad\text{with }\mathcal{N}_{s,l}=(2s+1)\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}$.
The fuzzy spherical harmonics $\hat{Y}_{l,m}$ satisfy the algebra Chu et al. (2001)
$\hat{Y}_{l,m}\hat{Y}_{l^{\prime},m^{\prime}}=\sum_{L,M}\,A^{l,l^{\prime},L}_{m,m^{\prime},M}\hat{Y}_{L,M},$ (29)
$\text{with }A^{l,l^{\prime},L,s}_{m,m^{\prime},M}=\sqrt{2s+1}(-1)^{2s+l+l^{\prime}+M}\sqrt{(2l+1)(2l^{\prime}+1)(2L+1)}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-M\end{pmatrix}\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix},$
where the coefficients in the expansion match the coefficients in the product of two spherical harmonics $Y_{l,m}$. Using the identity
$\left[c^{\dagger}_{m_{1},\sigma}c_{m_{2},\lambda},c^{\dagger}_{m^{\prime}_{1},\sigma^{\prime}}c_{m^{\prime}_{2},\lambda^{\prime}},\right]=\delta_{\lambda\sigma^{\prime}}\delta_{m_{2},m^{\prime}_{1}}c^{\dagger}_{m_{1},\sigma}c_{m^{\prime}_{2},\lambda^{\prime}}-\delta_{\sigma\lambda^{\prime}}\delta_{m_{1},m^{\prime}_{2}}c^{\dagger}_{m^{\prime}_{1},\sigma^{\prime}}c_{m_{2},\lambda},$
the commutator of the $n^{A}_{l,m}$ density modes can be expanded as
$\left[n^{A}_{l,m},n^{B}_{l^{\prime},m^{\prime}}\right]=\mathcal{N}_{s,l}\mathcal{N}_{s,l^{\prime}}\Bigg($ $\sum_{m_{1},m_{2},m^{\prime}_{2}}$ $(\hat{Y}_{l,m})_{m_{1},m_{2}}(\hat{Y}_{l^{\prime},m^{\prime}})_{m_{2},m^{\prime}_{2}}c^{\dagger}_{m_{1},\sigma}(A\cdot B)_{\sigma\lambda^{\prime}}c_{m^{\prime}_{2},\lambda^{\prime}}$
- $\sum_{m_{1},m_{2},m^{\prime}_{1}}$ $(\hat{Y}_{l^{\prime},m^{\prime}})_{m^{\prime}_{1},m_{1}}(\hat{Y}_{l,m})_{m_{1},m_{2}}c^{\dagger}_{m^{\prime}_{1},\sigma^{\prime}}(B\cdot A)_{\sigma^{\prime}\lambda}c_{m_{2},\lambda}\Bigg).$
Splitting $A\cdot B$ and $B\cdot A$ into commutator and anti-commutator and relabeling dummy indices gives
$\left[n^{A}_{l,m},n^{B}_{l^{\prime},m^{\prime}}\right]=\mathcal{N}_{s,l}\mathcal{N}_{s,l^{\prime}}\Bigg(\sum_{m_{1},m_{2}}$ $\left[\hat{Y}_{l,m},\hat{Y}_{l^{\prime},m^{\prime}}\right]_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}\{A,B\}_{\sigma\lambda}c_{m_{2},\lambda}$ (30)
$+\left\{\hat{Y}_{l,m},\hat{Y}_{l^{\prime},m^{\prime}}\right\}_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}[A,B]_{\sigma\lambda}c_{m_{2},\lambda}\Bigg)$
Now note that in the product of the $\hat{Y}_{l,m}$ matrices in (29), the 6j symbol is invariant under exchanging $l\to l^{\prime}$ while the 3j symbol picks up a $(-1)^{l+l^{\prime}+L}$ sign. Therefore,
$\left[n^{A}_{l,m},n^{B}_{l^{\prime},m^{\prime}}\right]=\mathcal{N}_{s,l}\mathcal{N}_{s,l^{\prime}}\sum_{m_{1},m_{2}}\sum_{L,M}\Bigg($ $\left(1-(-1)^{l+l^{\prime}+L}\right)A^{l,l^{\prime},L,s}_{m,m^{\prime},M}(\hat{Y}_{L,M})_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}\{A,B\}_{\sigma\lambda}c_{m_{2},\lambda}$
$+\left(1+(-1)^{l+l^{\prime}+L}\right)A^{l,l^{\prime},L,s}_{m,m^{\prime},M}(\hat{Y}_{L,M})_{m_{1},m_{2}}c^{\dagger}_{m_{1},\sigma}[A,B]_{\sigma\lambda}c_{m_{2},\lambda}\Bigg)$
The $\{A,B\}$ anticommutator only has support when $l+l^{\prime}+L$ is odd, while the $[A,B]$ commutator only has support when $l+l^{\prime}+L$ is even. After inserting the expressions for $\mathcal{N}_{s,l}$ and $A^{l,l^{\prime},L,s}_{m,m^{\prime},M}$, this reduces to the commutator (9).
Appendix C Central extensions
The planar GMP algebra admits a central extension, giving rise to the W-infinity algebra Cappelli and Maffi (2021). The Lie algebra of our density modes $\{n^{A}_{l,m}\}_{l\geq 1}$ at fixed $s$ is finite-dimensional; assuming it is semisimple as well, Whitehead’s lemma implies that any putative central extension is cohomologically trivial. Nevertheless, it is still natural to search for central terms at finite $s$ and ask whether they can survive in a suitable large-$s$ limit, potentially matching the central extension of area-preserving diffeomorphisms of $S^{2}$ discussed in Janssens and Wang (2026).
Ignoring the spin dof (i.e. setting $A=\text{id}$ in all $n^{A}_{l,m}$), a central extension should take the form
$\left[n_{l,m},n_{l^{\prime},m^{\prime}}\right]=-2\hskip-8.53581pt\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ odd}\end{subarray}}\hskip-5.69054pth_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n_{L,m+m^{\prime}}+\delta_{m+m^{\prime},0}\;g(s,l,l^{\prime},m)\mathbb{I}.$ (31)
The $\delta_{m+m^{\prime},0}$ delta function is necessary because of the grading of the algebra: $n_{1,0}$ acts as a grading operator, $n_{1,0},n_{l,m}]\propto m\,n_{l,m}$ (similar to $L_{0}$ in Virasoro). Consistency of the commutator with this grading requires the central extension term to vanish unless $m+m^{\prime}=0$.
The function $g$ must satisfy $g(s,l,l^{\prime},m)=-g(s,l^{\prime},l,-m)$ so that $n_{l,m},n_{l^{\prime},m^{\prime}}]=-[n_{l^{\prime},m^{\prime}},n_{l,m}]$ is not violated. Further constraints come from the Jacobi identity:
$\sum_{\begin{subarray}{c}L=0\\ l_{1}+l_{2}+L\text{ odd}\end{subarray}}h_{L}(s,l_{1},l_{2})(-1)^{m_{1}+m_{2}}\begin{pmatrix}l_{1}&l_{2}&L\\ m_{1}&m_{2}&m_{3}\end{pmatrix}g(s,l_{3},L,m_{3})+\text{ cyclic permutations }=0,$
assuming $m_{1}+m_{2}+m_{3}=0$. One possible solution that satisfies the Jacobi identity is
$g(s,l,l^{\prime},m)=\delta_{l,l^{\prime}}(-1)^{m}\,m\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}^{2}.$
We leave it to future work to determine whether this solution is connected to the central extension of the planar GMP algebra in the $s\to\infty$ limit with $l\sim\sqrt{s}$, or related to other central extensions for 3d CFTs Janssens and Wang (2026).
Appendix D No $so(5)$ representation extending $so(3)$ representation generators for $s>\tfrac{1}{2}$
In this appendix we consider the general class of fermion-bilinear operators $n^{A}(X)$ and ask whether one can extend the orbital rotation generators $J_{a}=n^{\mathbb{I}}(L_{a})$ to a larger $so(5)$ (or $so(3,2)$) Lie algebra. We show that for $s>\tfrac{1}{2}$ this is not possible, independent of the internal spin space dimension.
A fermion bilinear built from $2s+1$ orbital modes and $d$ internal spin/flavor degrees of freedom can be written as
$n^{A}(X):=\sum_{m_{1},m_{2}=-s}^{s}\sum_{\sigma,\lambda=1}^{d}X_{m_{1}m_{2}}\,c^{\dagger}_{m_{1},\sigma}\,A_{\sigma\lambda}\,c_{m_{2},\lambda},$
where $X$ is a $(2s+1)\times(2s+1)$ matrix acting on the orbital indices $m$, and $A$ is a $d\times d$ matrix acting on the internal indices $\sigma$. The density modes $n^{A}_{l,m}$ in (11) are the special case $X=\hat{Y}_{l,m}$ (the fuzzy spherical harmonics) with $d=2$ corresponding to spin $\uparrow,\downarrow$. Using the canonical anticommutation relations for the fermions, the commutator of two such bilinears closes:
$\bigl[n^{A}(X),\,n^{B}(Y)\bigr]$ $=n^{AB}(XY)-n^{BA}(YX)=\frac{1}{2}\,n^{\{A,B\}}\!\bigl([X,Y]\bigr)+\frac{1}{2}\,n^{[A,B]}\!\bigl(\{X,Y\}\bigr).$
The corresponding special case for $n^{A}_{l,m}$ is (30).
The $so(3)$ rotation generators are
$J_{z}=n^{\mathbb{I}}(L_{z}),\hskip 18.49988ptJ_{\pm}=n^{\mathbb{I}}(L_{\pm}),$
where $\mathbb{I}$ denotes the $d\times d$ identity and $L_{z},L_{\pm}$ are the generators of the spin-$s$ irrep of $so(3)$. We want to find an additional fermion-bilinear generator $D_{z}=n^{A_{z}}(X_{z})$ that commutes with all rotations. Since $L_{a},D_{z}]=0$ implies $L_{a},X_{z}]=0$ for $a=z,\pm$, Schur’s lemma for the irreducible spin-$s$ orbital representation gives $X_{z}\propto\mathbb{I}_{2s+1}$, so we may take
$D_{z}=n^{A_{z}}(\mathbb{I})=\sum_{m_{1}=-s}^{s}c^{\dagger}_{m_{1},\sigma}(A_{z})_{\sigma\lambda}c_{m_{1},\lambda}.$
In other words, rotational invariance restricts $D_{z}$ to an onsite term that acts independently on each orbital $m$.
Next, impose $J_{z},D_{\pm}]=0$. Writing $D_{\pm}=n^{A_{\pm}}(X_{\pm})$, the commutator formula reduces to
$[J_{z},D_{\pm}]=n^{A_{\pm}}([L_{z},X_{\pm}]),$
[…]
$[D_{+},D_{-}]=\frac{1}{2}\,n^{[A_{+},A_{-}]}(\{X_{+},X_{-}\}).$
Requiring $D_{+},D_{-}]$ to be proportional to $D_{z}=n^{A_{z}}(\mathbb{I})$ implies
$[A_{+},A_{-}]\propto A_{z},\hskip 18.49988pt\{X_{+},X_{-}\}\propto\mathbb{I}_{2s+1}.$
If the Fock-space inner product is to be invariant for the $so(5)$ representation, we can choose conventions such that $D_{-}=(D_{+})^{\dagger}$. Since $(n^{A}(X))^{\dagger}=n^{A^{\dagger}}(X^{\dagger})$, this gives $A_{-}=A_{+}^{\dagger}$ and $X_{-}=X_{+}^{\dagger}$. Writing $X_{+}=\mathrm{diag}(k_{-s},\dots,k_{s})$, the condition $\{X_{+},X_{-}\}\propto\mathbb{I}$ becomes
$\{X_{+},X_{+}^{\dagger}\}=2\,\mathrm{diag}(|k_{m}|^{2})\propto\mathbb{I}\qquad\Longrightarrow\qquad|k_{m}|=\text{const}.$
With the normalization used above, this is $|k_{m}|=1/\sqrt{2}$, i.e. $k_{m}=e^{i\theta_{m}}/\sqrt{2}$.
A further relation required by the $so(5)$ algebra in this ladder basis is
$[J_{-},[J_{-},D_{+}]]=0.$
Using the closure of bilinears and that $J_{-}=n^{\mathbb{I}}(L_{-})$, this reduces to
$[J_{-},[J_{-},D_{+}]]=n^{A_{+}}\!\bigl([L_{-},[L_{-},X_{+}]]\bigr),$
[…]
For $s>\tfrac{1}{2}$, the coefficients $c_{m}c_{m-1}$ are nonzero for $m=-s+2,\dots,s$, hence
$k_{m}-2k_{m-1}+k_{m-2}=0\hskip 18.49988pt(m=-s+2,\dots,s).$
Thus $k_{m}$ has vanishing second finite difference and must be affine:
$k_{m}=am+b.$
Combining this with $|k_{m}|=1/\sqrt{2}$ for all $m$ forces $a=0$ whenever there are at least three distinct $m$-values, i.e. for $s\geq 1$. Hence $k_{m}$ must be constant, so $X_{+}\propto\mathbb{I}$ and therefore $D_{+}$ commutes with all $J_{a}$, contradicting the requirement that $D_{\pm}$ provide a nontrivial extension of the rotation algebra. The case $s=\tfrac{1}{2}$ evades this obstruction because $(L_{-})^{2}=0$, so $L_{-},[L_{-},X_{+}]]=0$ holds identically.
For $s>\tfrac{1}{2}$, there is no choice of diagonal $X_{+}$ with $|k_{m}|=\mathrm{const}$ that satisfies the mixed relation $J_{-},[J_{-},D_{+}]]=0$ while keeping $D_{+}$ nontrivial. This obstruction is independent of the internal dimension $d$. In particular, there is no $so(5)$ representation on these fermion bilinears (and most likely also not $so(3,2)$ one) that extends the spin-$s$ orbital $so(3)$ generators and is unitary with respect to the Fock inner product for $s>\tfrac{1}{2}$.
Appendix E Larger $so(3,2)$ representations from the Lie algebra coproduct
As a complementary construction, one may enlarge the minimal $s=\tfrac{1}{2}$ representation by using the standard coproduct on the universal enveloping algebra $U(so(3,2))$. For any generator $x\in so(3,2)$ we define
$\Delta(x):=x\otimes\mathbb{I}+\mathbb{I}\otimes x,$
and obtain an $n$-fold coproduct $\Delta^{(n)}$ by iteration.
While this produces an $so(3,2)$ representation on a larger Hilbert space, it does not match the notion of “increasing the fuzzy-sphere size” relevant for the critical models. In particular, the induced $so(3)$ rotations act on the tensor product Hilbert space $\mathcal{H}_{s=\frac{1}{2}}^{\otimes n}$ and therefore furnish the reducible representation $(V^{\frac{1}{2}})^{\otimes n}$ rather than a single spin-$s$ irrep $V^{s}$ with $s=\tfrac{n}{2}$. Consequently, the rotation generators obtained from $\Delta^{(n)}$ are not proportional to the $l=1$ density modes $n^{0,s}_{1m}$ of a spin-$s$ fuzzy sphere.
One may also deform the coproduct by a Drinfeld twist. For an invertible $F\in H\otimes H$ with $H=U(so(3,2))$, the twisted coproduct is
$\Delta^{F}(x):=F\,\Delta(x)\,F^{-1},$
which still satisfies $\Delta^{F}(x),\Delta^{F}(y)]=\Delta^{F}([x,y])\,$. However, since $\Delta^{F}(x)$ is conjugate to $\Delta(x)$, twisting does not change the equivalence class of the resulting tensor-product representation. In particular, it cannot convert the coproduct rotations on $(V^{\frac{1}{2}})^{\otimes n}$ into the single-irrep rotations appropriate to a spin-$s$ fuzzy sphere, and therefore does not resolve the structural mismatch with the thermodynamic limit of the critical fuzzy-sphere models.

### v2

@@
 The fuzzy sphere geometry admits two thermodynamic limits: a local planar limit yielding the fuzzy plane, and a commutative limit yielding an ordinary sphere.
-In the planar limit, high-angular-momentum modes recover the Girvin–MacDonald–Platzman algebra, whereas in the commutative limit the low-angular-momentum modes become semiclassical.
-We further find an explicit representation of the conformal algebra $so(3,2)$ in the minimal two-electron system and extend it to larger systems via an $so(3)$ equivariant coproduct.
-Because the coproduct splits one $so(3)$ representation into a tensor product, it is structurally mismatched with the thermodynamic limit of critical fuzzy sphere Hamiltonians.
+In the planar limit, high-angular-momentum modes recover the Girvin–MacDonald–Platzman algebra, whereas in the commutative limit, the low-angular-momentum modes become semiclassical.
+Upon further restricting to a subspace with few spin flips above the paramagnetic reference state, they behave approximately as harmonic oscillators.
+We also find an explicit representation of the conformal algebra $so(3,2)$ in the minimal two-electron system and extend it to larger systems via an $so(3)$ equivariant coproduct.
+Because the coproduct splits one $so(3)$ representation into a tensor product, it is structurally mismatched with the thermodynamic limit of critical fuzzy sphere models.
 1 Introduction
@@
 The fuzzy sphere model describes spinful electrons on the 2-sphere with a monopole at the origin, projected to the lowest Landau level (LLL).
-Remarkably, it reproduces scaling dimensions of the 3d Ising CFT with good precision already in small systems of only $\mathcal{O}(10)$ electrons.
+Within a two-parameter family of $SO(3)$ invariant Hamiltonians, one can tune to a critical point whose low-lying spectrum matches 3d Ising CFT scaling dimensions with good precision already for $\mathcal{O}(10)$ electrons.
+We emphasize, however, that obtaining CFT-like spectra from small 2+1d systems is not unique to the fuzzy sphere: similar behavior occurs in more conventional discretizations, such as the 2+1d transverse-field Ising model on the icosahedron Lao and Rychkov (2023), where good accuracy requires conformal perturbation theory.
+Conformal perturbation theory is likewise useful on the fuzzy sphere to further improve the extracted scaling dimensions Läuchli et al.
+(2025); Voinea et al.
+(2025).
+A central conceptual problem is locating the conformal points of the Hamiltonian.
+Possible routes include conformal perturbation theory and enlarged symmetry, but we leave this challenging question to the future.
 The Landau-level–projected density operators that enter the fuzzy-sphere Hamiltonians generate an intriguing algebra that is of independent interest, and which we study in this paper.
@@
 The commutative, semiclassical regime is expected to be the more relevant one for the low-energy spectrum of the critical fuzzy-sphere Hamiltonians.
+We also refine the observation of He (2025) that, after restricting to a low-excitation subspace above the paramagnetic state, the density modes admit an approximate harmonic-oscillator description.
 In 2+1 dimensions, conformal transformations are described by the finite-dimensional Lie algebra $so(3,2)$, reviewed in Section 2.
@@
 $\{\psi_{\sigma}(\Omega),\psi^{\dagger}_{\sigma}(\Omega^{\prime})\}=K(\Omega,\Omega^{\prime}),\;\text{ where}$ (4)
-$K(\Omega,\Omega^{\prime})=\frac{1}{R}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\left(Y^{(s)}_{s,m}(\Omega^{\prime})\right)^{*}=\frac{2s+1}{R^{2}}e^{2is\alpha(\Omega,\Omega^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\text{ and }$
+$K(\Omega,\Omega^{\prime})=\frac{1}{R^{2}}\sum_{m=-s}^{s}Y^{(s)}_{s,m}(\Omega)\left(Y^{(s)}_{s,m}(\Omega^{\prime})\right)^{*}=\frac{2s+1}{R^{2}}e^{2is\alpha(\Omega,\Omega^{\prime})}\cos^{2s}{\left(\frac{\gamma}{2}\right)},\text{ and }$
 […]
@@
 (2025), and the free fermion Zhou et al.
-(2025a) and Chern-Simons-matter CFTs Zhou and He (2025); Zhou et al.
+(2025a); Voinea et al.
+(2026) and Chern-Simons-matter CFTs Zhou and He (2025); Zhou et al.
 (2025b).
+In addition, fractional-filling fuzzy-sphere models (rather than integer $\nu=1$ filling) can also realize 3d CFTs Voinea et al.
+(2025).
 3.3 Conformal generators for the fuzzy sphere Hamiltonian
@@
 Integrating over $\Omega$ gives the corresponding quadratic relations among density modes:
-$\sum_{l,m}(-1)^{m}n^{\alpha}_{l,m}\,n^{\alpha}_{l,-m}=2n^{0}_{0,0}-\sum_{l,m}(-1)^{m}n^{0}_{l,m}\,n^{0}_{l,-m},\hskip 18.49988pt\alpha=x,y,z.$ (13)
+$\sum_{l,m}(-1)^{m}n^{\alpha}_{l,m}\,n^{\alpha}_{l,-m}=2(2s+1)n^{0}_{0,0}-\sum_{l,m}(-1)^{m}n^{0}_{l,m}\,n^{0}_{l,-m},\hskip 18.49988pt\alpha=x,y,z.$ (13)
 4.2 Jacobi identity and commuting subalgebras
 Does the mode algebra, viewed abstractly through the commutation relations (9), define a genuine Lie algebra?
-Let $\mathcal{A}_{s}=\{\tilde{n}^{A}_{l,m}\}_{l=0,|m|\leq l}$ be the algebra spanned by formal generators $\tilde{n}^{A}_{l,m}$, equipped with the bracket defined by (9).
+Let $\mathcal{A}_{s}=\{\tilde{n}^{A}_{l,m}\}_{0\leq l=0\leq 2s,|m|\leq l}$ be the algebra spanned by formal generators $\tilde{n}^{A}_{l,m}$, equipped with the bracket defined by (9).
 A priori this bracket need not satisfy Jacobi.
@@
 For the full algebra generated by the $\tilde{n}^{A}_{l,m}$, the rank doubles to $2N$, and a natural maximal commuting subalgebra is generated by $\{\tilde{n}^{0}_{l,0},\,\tilde{n}^{z}_{l,0}\}$.
-Finally, a separate question is whether this algebra admits a nontrivial central extension; we briefly comment on this in Appendix C.
+Finally, a separate question is whether this algebra admits a nontrivial central extension; we briefly comment on this in Appendix D.
 5 Thermodynamic limit of the density mode algebra
@@
 For the critical fuzzy-sphere Hamiltonians, we expect the low-energy sector to be controlled by this latter, semiclassical regime.
+Moreover, Sec.
+5.3 shows that if one additionally focuses on a low-excitation sector consisting of a few low-momentum spin flips above the paramagnetic reference state, the density modes admit an approximate harmonic-oscillator description.
 5.1 Quantum planar limit of large-$l$ density modes
@@
 The operators $e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}$ satisfy the algebra following from (19),
-$e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}\,e^{i\mathbf{p}\cdot\hat{\mathbf{x}}}=\exp\!\Big(-\frac{i}{2}\Theta\,(\mathbf{q}\times\mathbf{p})\Big)\,e^{i(\mathbf{q}+\mathbf{p})\cdot\hat{\mathbf{x}}},\hskip 18.49988pt\mathbf{q}\times\mathbf{p}:=q_{1}p_{2}-q_{2}p_{1}.$ (22)
+$e^{i\mathbf{q}\cdot\hat{\mathbf{x}}}\,e^{i\mathbf{p}\cdot\hat{\mathbf{x}}}=\exp\!\Big(-\frac{i}{2}\Theta\,(\mathbf{q}\times\mathbf{p})\Big)\,e^{i(\mathbf{q}+\mathbf{p})\cdot\hat{\mathbf{x}}},\hskip 18.49988pt\mathbf{q}\times\mathbf{p}:=q_{1}p_{2}-q_{2}p_{1}.$
 Using the fermion bilinear expansion (21), this implies the spin-enriched planar GMP algebra
@@
 We can use the large-$l$ correspondence (20) to rewrite the fuzzy-sphere Hamiltonian (7) in the planar limit. Substituting $n^{A}_{l,m}\mapsto\rho^{A}_{\mathbf{q}}$ (with $|\mathbf{q}|=l/R$) yields
-$H\approx\rho^{x}_{\mathbf{0}}+\int d^{2}q\;V(q)\Big(\rho^{z}_{\mathbf{q}}\rho^{z}_{-\mathbf{q}}+\rho^{0}_{\mathbf{q}}\rho^{0}_{-\mathbf{q}}\Big),\hskip 18.49988ptV(q)=U_{0}+U_{1}q^{2}.$ (26)
+$H\approx-h\rho^{x}_{\mathbf{0}}+\int d^{2}q\;V(q)\Big(\rho^{0}_{\mathbf{q}}\rho^{0}_{-\mathbf{q}}-\rho^{z}_{\mathbf{q}}\rho^{z}_{-\mathbf{q}}\Big),\hskip 18.49988ptV(q)=U_{0}+U_{1}q^{2}.$ (25)
 The first term acts as a uniform field, and the remaining terms describe planar density–density interactions.
@@
 However, this naive semiclassical bracket does not directly reproduce several relations that appear to emerge in the low-energy sector of critical fuzzy-sphere Hamiltonians.
-For instance, Ref.
-He (2025) found that, upon restricting to states near the paramagnetic ground state, the operators $n^{+}_{l,m}$ and $n^{-}_{l,m}$ approximately behave as oscillator modes, with $n^{+}_{l,m},n^{-}_{l^{\prime},m^{\prime}}]\approx\delta_{l,l^{\prime}}\delta_{m+m^{\prime},0}n^{z}_{0,0}$.
-By contrast, the semiclassical commutator (28) generically yields a sum over many $L$ rather than collapsing to a single $L=0$ term.
-Similarly, the semiclassical commutator (27) does not by itself imply the conformal commutation relations observed numerically in the same low-energy subspace, e.g.
+For example, the semiclassical commutator (26) does not by itself imply the conformal commutation relations observed numerically in the same low-energy subspace, e.g.
 $\Lambda_{z},[\Lambda_{z},D]]=4D$ in Fardelli et al.
@@
 2.4.
+5.3 Oscillator approximation in subspace above paramagnetic state
+[…]
+The commutators $n^{+}_{l,m},n^{+}_{l^{\prime},m^{\prime}}]=[n^{-}_{l,m},n^{-}_{l^{\prime},m^{\prime}}]=0$ hold exactly, while the mixed commutator is given by
+$\left[n^{+}_{l,m},n^{-}_{l^{\prime},m^{\prime}}\right]=$ $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ even}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}$ (28)
+- $\sum_{\begin{subarray}{c}L=0\\ l+l^{\prime}+L\text{ odd}\end{subarray}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{0}_{L,m+m^{\prime}}.$
+Following He (2025), we consider the subspace spanned by states with a small number $k$ of spin flips and bounded angular momentum,
+[…]
+On this subspace, the density mode $n^{z}_{0,0}$ is extensive, with eigenvalue N+2k-N+2k on a state with $k$ spin flips.
+In contrast, for $l>0$ the density modes satisfy the operator norm bound (see Appendix C for derivation)
+$\|n^{z}_{l,m}\|_{P_{k}}\leq 2k\sqrt{2l+1},$
+where the $P_{k}$ subscript denotes the restriction to the sector with $k$ up spins. The same bound holds for $n^{0}_{l,m}$.
+Although (28) contains many $n^{z}_{L,m}$ terms, on the subspace of $k$ spin flips with $l_{i}\leq l_{\text{max}}$ only those with $L\leq 2kl_{\text{max}}$ survive.
+The reason is that $n^{z}_{L,m}$ carries angular momentum $L$, so for $L>2kl_{\text{max}}$ it necessarily takes any state inside the subspace (29) to states with angular momentum greater than $kl_{\text{max}}$ outside the subspace.
+Therefore, the operator norm of the remaining contribution can be bounded as
+$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
+[…]
+for some constants $C$, $c$ independent of $s$. The inequality is derived in Appendix C.
+In particular, if $l,l^{\prime}\leq l_{\text{max}}$ and $kl_{\text{max}}=o(\sqrt{s})$, then the right-hand side divided by $N$ vanishes as $s\to\infty$.
+Then the rescaled modes satisfy the oscillator approximation on this subspace,
+$\Bigg[\frac{n^{+}_{l,m}}{\sqrt{N}},\,\frac{n^{-}_{l^{\prime},m^{\prime}}}{\sqrt{N}}\Bigg]\;\approx\;\delta_{l,l^{\prime}}\,\delta_{m+m^{\prime},0}\,\frac{n^{z}_{0,0}}{N}\;\approx\;\delta_{l,l^{\prime}}\,\delta_{m+m^{\prime},0}.$ (30)
+Sharpening the bounds on the $3j$ and $6j$ symbols used in the derivation in Appendix C would likely improve the dependence on $k$ and $l_{\text{max}}$, but we do not pursue this here.
 6 Hopf algebra construction of $so(3,2)$ representations
@@
 The $R^{so(5)}_{\pm}$ and $S^{so(5)}_{\pm}$ operators are then defined by the same commutation relations as before, but with $D_{\pm}$ replaced with $D_{\pm}^{so(5)}$.
-In Appendix D, we show that this so(3,2)so(3,2)/$so(5)$ representation built from the $n^{A}_{l,m}$ density modes at $s=\tfrac{1}{2}$ does not extend to larger $s$, even if we allow an arbitrary enlargement of the internal spin degrees of freedom.
+In Appendix E, we show that this so(3,2)so(3,2)/$so(5)$ representation built from the $n^{A}_{l,m}$ density modes at $s=\tfrac{1}{2}$ does not extend to larger $s$, even if we allow an arbitrary enlargement of the internal spin degrees of freedom.
 For example, it is easy to see that the commutator $D_{+},D_{-}]$ produces an additional $n^{z}_{2,0}$ term when $s>\tfrac{1}{2}$.
@@
 While this coproduct gives us a $so(3,2)$ representation defined on a larger Hilbert space, it is different from the thermodynamic limit of the critical fuzzy sphere models, which takes $s\to\infty$ within a single $V^{s}$.
-Using the standard coproduct of Lie algebras leads to a similar mismatch, see Appendix E.
+Using the standard coproduct of Lie algebras leads to a similar mismatch, see Appendix F.
 Nevertheless, it seems interesting that this coproduct induces a nontrivial $so(3,2)$ representation on a tensor product Hilbert space.
@@
 After inserting the expressions for $\mathcal{N}_{s,l}$ and $A^{l,l^{\prime},L,s}_{m,m^{\prime},M}$, this reduces to the commutator (9).
-Appendix C Central extensions
+Appendix C Derivation of oscillator approximation
+[…]
+In the $k=1$ subspace spanned by states $|\ket{a;b}:=c^{\dagger}_{a,\uparrow}c_{b,\downarrow}\ket{\Downarrow}$, the matrix elements of $n^{z}_{l,m}$ with $l>0$ are
+[…]
+$=\mathcal{N}_{s,l}\left((\hat{Y}_{l,m})_{a^{\prime},a}\delta_{b,b^{\prime}}+(\hat{Y}_{l,m})_{b,b^{\prime}}\delta_{a,a^{\prime}}\right)$
+Then using the triangle inequality for the two terms, we can bound the operator norm of $n^{z}_{l,m}$ with $l>0$ in this subspace by
+$\|n^{z}_{l,m}\|\leq 2\mathcal{N}_{s,l}\|\hat{Y}_{l,m}\|.$
+To bound the operator norm of $\hat{Y}_{l,m}$, we use that the norm is equal to the square root of the largest eigenvalue of $\hat{Y}^{\dagger}_{l,m}\hat{Y}_{l,m}$. This is a diagonal matrix,
+$(\hat{Y}^{\dagger}_{l,m}\hat{Y}_{l,m})_{bb^{\prime}}=\sum_{a}(\hat{Y}_{l,m})^{*}_{ab}(\hat{Y}_{l,m})_{ab^{\prime}}=\delta_{b,b^{\prime}}\sum_{a}|(\hat{Y}_{l,m})_{ab}|^{2}=\delta_{b,b^{\prime}}|(\hat{Y}_{l,m})_{b+m,b}|^{2},$
+because $(\hat{Y}_{l,m})_{ab}$ is nonzero only when b=mab=m-a and $(\hat{Y}_{l,m})_{ab^{\prime}}$ is nonzero only when $b^{\prime}=m-a$. Therefore, the operator norm of $\hat{Y}_{l,m}$ is equal to
+$\|\hat{Y}_{l,m}\|=\max_{b}|(\hat{Y}_{l,m})_{b+m,b}|=\sqrt{2l+1}\max_{b}\left|\begin{pmatrix}s&s&l\\ b+m&-b&-m\end{pmatrix}\right|\leq\sqrt{\frac{2l+1}{2s+1}}.$
+The last inequality follows from the condition
+$(2j_{3}+1)\sum_{m_{1},m_{2}}\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}^{2}=\begin{cases}1&|j_{1}-j_{2}|\leq j_{3}\leq j_{1}+j_{2}\\ 0&\text{else }\end{cases},$
+which implies
+$\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}^{2}\leq\frac{1}{2\max{\{j_{1},j_{2},j_{3}\}}+1}.$
+Together with the fact that $\mathcal{N}_{s,l}=(2s+1)\begin{pmatrix}s&s&l\\ s&-s&0\end{pmatrix}\leq\sqrt{2s+1}$, this leads to
+$\|n^{z}_{l,m}\|\leq 2\sqrt{2l+1}.$
+In contrast, for $n^{z}_{0,0}$ we have
+[…]
+therefore $n^{z}_{0,0}$ is extensive in $N$ while $\|n^{z}_{l,m}\|=\mathcal{O}(\sqrt{l})$ for $l>0$ (and analogously for $n^{0}_{l,m}$).
+In the sector with $k$ spin flips spanned by states $|\ket{A;B}:=c^{\dagger}_{a_{1}}\dots c^{\dagger}_{a_{k}}c_{b_{1}}\dots c_{b_{k}}\ket{\Downarrow}$ with $A=\{a_{1},\dots,a_{k}\}$ and $B=\{b_{1},\dots,b_{k}\}$, the matrix elements of $n^{z}_{l,m}$ are
+[…]
+Using again the triangular inequality for the $2k$ terms, the operator norm is then bounded as
+$\|n^{z}_{l,m}\|\leq 2k\mathcal{N}_{s,l}\|\hat{Y}_{l,m}\|\leq 2k\sqrt{2l+1}.$
+Again, the same holds for $n^{0}_{l,m}$. In contrast, the operator norm of $n^{z}_{0,0}$ is again extensive, concretely
+[…]
+[…]
+$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
+$\leq\left(\sum_{L=1}^{2kl_{\text{max}}}(h_{L}(s,l,l^{\prime}))^{2}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}^{2}\right)^{1/2}\left(\sum_{L=1}^{2kl_{\text{max}}}\|n^{z}_{L,m+m^{\prime}}\|^{2}_{P_{k}}\right)^{1/2}$
+To bound the first factor, we use $\begin{pmatrix}j_{1}&j_{2}&j_{3}\\ m_{1}&m_{2}&m_{3}\end{pmatrix}\leq\frac{1}{\sqrt{2\max{\{j_{1},j_{2},j_{3}\}}+1}}$ on three different 3j symbols,
+$\sum_{L=1}^{2kl_{\text{max}}}(h_{L}(s,l,l^{\prime}))^{2}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}^{2}\leq(2l+1)(2l^{\prime}+1)\sum_{L=1}^{2kl_{\text{max}}}\frac{\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix}^{2}}{\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{2}}.$
+Orthogonality of the 6j symbols then gives
+$\sum_{L=1}^{2kl_{\text{max}}}\frac{\begin{Bmatrix}l&l^{\prime}&L\\ s&s&s\end{Bmatrix}^{2}}{\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{2}}\leq\frac{1}{2s+1}\max_{L=1,\dots,2kl_{\text{max}}}\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{-2}$
+The inverse squared 3j symbol is known to be
+$\begin{pmatrix}s&s&L\\ s&-s&0\end{pmatrix}^{-2}=\frac{(2s-L)!(2s+L+1)!}{((2s)!)^{2}}.$
+For $L_{\text{max}}=2kl_{\text{max}}\leq s/2$, this can be bounded using Stirling’s approximation as
+[…]
+where $c$ and $C$ are some constants. If $L_{\text{max}}\ll\sqrt{s}$, this exponential is $\mathcal{O}(1)$. Altogether, this gives the bound
+$\left\|\sum_{L=1}^{2kl_{\text{max}}}h_{L}(s,l,l^{\prime})(-1)^{m+m^{\prime}}\begin{pmatrix}l&l^{\prime}&L\\ m&m^{\prime}&-m-m^{\prime}\end{pmatrix}n^{z}_{L,m+m^{\prime}}\right\|_{P_{k}}$
+[…]
+Appendix D Central extensions
 The planar GMP algebra admits a central extension, giving rise to the W-infinity algebra Cappelli and Maffi (2021).
@@
 We leave it to future work to determine whether this solution is connected to the central extension of the planar GMP algebra in the $s\to\infty$ limit with $l\sim\sqrt{s}$, or related to other central extensions for 3d CFTs Janssens and Wang (2026).
-Appendix D No $so(5)$ representation extending $so(3)$ representation generators for $s>\tfrac{1}{2}$
+Appendix E No $so(5)$ representation extending $so(3)$ representation generators for $s>\tfrac{1}{2}$
 In this appendix we consider the general class of fermion-bilinear operators $n^{A}(X)$ and ask whether one can extend the orbital rotation generators $J_{a}=n^{\mathbb{I}}(L_{a})$ to a larger $so(5)$ (or $so(3,2)$) Lie algebra.
@@
 In particular, there is no $so(5)$ representation on these fermion bilinears (and most likely also not $so(3,2)$ one) that extends the spin-$s$ orbital $so(3)$ generators and is unitary with respect to the Fock inner product for $s>\tfrac{1}{2}$.
-Appendix E Larger $so(3,2)$ representations from the Lie algebra coproduct
+Appendix F Larger $so(3,2)$ representations from the Lie algebra coproduct
 As a complementary construction, one may enlarge the minimal $s=\tfrac{1}{2}$ representation by using the standard coproduct on the universal enveloping algebra $U(so(3,2))$.